A body moves along the x-axis. Its position is expressed in the given equation with respect to time: x = -4t + 2t2

. With the given time intervals of t = 0 to t = 1 s and t = 1 s to t = 3 s, find the body’s a) displacement b) average velocity and c) instantaneous velocity at t = 

2.5 s. 

1
Expert's answer
2021-04-29T07:27:43-0400

1. Let's consider the interval "[0s, 1s]". At the beginning of this interval the position of the object is:



"x_1 = -4\cdot 0s + 2\cdot (0s)^2 = 0"

At the end:



"x_2 = -4\cdot 1s + 2\cdot (1s)^2 = -2"

Thus, the displacement is:



"d_1 = x_2 - x_1 = -2"

By definition, the averabe velocity is:



"v_1 = \dfrac{d_1}{1s - 0s} = \dfrac{-2}{1s - 0s} = -2"

2. Let's consider the interval "[1s, 3s]". At the beginning of this interval the position of the object is "x_2 = -2" , at the end:



"x_3 = -4\cdot 3s + 2\cdot (3s)^2 = 6"

Thus, the displacement is:



"d_2 = x_3 - x_2 = 8"

By definition, the averabe velocity is:



"v_2 = \dfrac{d_2}{3s - 1s} = \dfrac{8}{3s - 1s} =4"

3. The instantaneous velocity is defined as the derivative of position with respect to time. Thus:



"v(t) = \dfrac{dx}{dt} = 4t-4\\
v(2.5s) = 4\cdot 2.5s - 4 = 6"

Answer. a) Displacements: -2 and 8, b) average velocities: -2 and 4, c) instantaneous velocity: 6.